OpenStudy (anonymous):

for f(x)=1/x-5 and g(x)=x^2 + 2 substitute the value of g (x) into the function f(x) in place of x find the value of f(g(x))

OpenStudy (anonymous):

Ok, so what's the problem? U see the x in the first equation, replace it by x^2 + 2 simplify.

OpenStudy (anonymous):

right so i got 1/x^2+2-5 can it be simplified farther than that?

OpenStudy (anonymous):

f(g(x)) = 1/(g(x) - 5) Now do the substitution for g(x) and simplify.

OpenStudy (anonymous):

wait is f(x) like this: \[f(x) = \frac{1}{x-5}\] Or like this: \[f(x) = \frac{1}{x}-5\]

OpenStudy (anonymous):

the first option

OpenStudy (anonymous):

Doesn't matter, whatever one, just replace the x.

OpenStudy (anonymous):

Well it does matter because one can be simplified and the other cannot ;p

OpenStudy (anonymous):

so is the answer 1/x^2-3?

OpenStudy (anonymous):

So 1/(x^2-3).

OpenStudy (anonymous):

So you said you got: \[f(g(x)) = \frac{1}{x^2 +2 - 5}\] Which I should hope you can see simplifies a little.

OpenStudy (anonymous):

Right. Please use parens if you aren't going to use the latex for your fraction what you wrote isn't the same as what the answer is. (what estudier wrote)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

polpak is right about being clear, better to use brackets..